1. Consider the following data for aluminum, Cu, and an unknown metal Z. metal Z 10. Metal Mass (8) Specific heat (J/8°C) Ini Just 1c, cant figure out how to calculate specific heat without the heat energy
Answer #1

Heat exchanged between 2 metals must be same so as the final temperature is acheived 48.5 C

Z initial temperature 55 C Al inital temperature 40 C

So Z releases heat and Al accepts it.

So for Al / Z combination Heat taken up by Al = m*CAl*\DeltaT = 10gm* (0.900 J/gm*C) * (48.5-40) C = 76.5 J

So Z takes up 76.5 J heat

We can write m*CZ*\DeltaT = 10gm* (Cz)* (55-48.5) C = 76.5 J

So C z = 1.17 J/gm*C